OpenStudy (anonymous):

What is the solution to the rational equation 3 over x - 2 = 2 over x -3 i just dont understand this one at all

OpenStudy (anonymous):

3 = 2 -- -- x-2 x-3 we can get rid of the fractions by multiply each term by the lcm of x-2 and x-3 which is (x-2)(x-3) 3(x-2)(x-3) 2(x-2)(x-3) -------- = --------- (x-2) x-3 3(x-3) = 2(x-2) 3x -9 = 2x - 4 the rest is easy

OpenStudy (anonymous):

add 3x to the other side to get x and then -4 to the other side -5 x=-5?

OpenStudy (anonymous):

positive 5

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you can solve any fractional equation by this method